OpenStudy (k8lyn911):

Calculus II - Average Values [0, pi] f(c) = f ave -> 10sin(c) - 5 sin(2c) = 20/pi Solve for c. (Use three decimal places. There should be two answers.)

myininaya (myininaya):

Was the initial question to solve 10sin(c)-5sin(2c)=10/pi ?

OpenStudy (k8lyn911):

The initial question was to find the average value of 10sin(x)-5sin(x) between o and pi. Now I need to find c such that f ave = f(c) rounded to three decimal places.

myininaya (myininaya):

another dumb question is it 10sin(x)-5sin(x) or 10sin(x)-5sin(2x)

OpenStudy (amistre64):

on an interval [a,b] of some function f the average is determined to be:\[\frac1{b-a}\int_{a}^{b}f(x)dx\] or does memory betray me?

myininaya (myininaya):

you are right

OpenStudy (k8lyn911):

That's right. It's 10sin(x)-5sin(2x).

OpenStudy (k8lyn911):

I already got the f ave part right, though, so I just need to solve for c.

OpenStudy (amistre64):

k :)

OpenStudy (amistre64):

sin(2c) = 2sin(c)cos(c) maybe?

OpenStudy (amistre64):

if so, factor the commons i believe

OpenStudy (k8lyn911):

Trig isn't my strong point. Could you walk me through it?

OpenStudy (amistre64):

an identity is: sin(2x) = 2sin(x)cos(x) so replacing it gets us: 10 sin(x) - 5(2sin(x)cos(x)) = K 10 sin(x) (1 - cos(x)) = K

OpenStudy (amistre64):

might help if we could get a ^2 going on tho

myininaya (myininaya):

maybe i'm giving up too soon but i'm not seeing a good algebraic/trig approach to solve this :(

OpenStudy (dan815):

show me the initial function you use to find the avg

OpenStudy (dan815):

"The initial question was to find the average value of 10sin(x)-5sin(x) between o and pi." you got x there is it suppsoed to be 2x?

OpenStudy (rational):

f(x) = 10sin(x) - 5sin(2x) Average value of f(x) in interval 0,pi is 20/pi

OpenStudy (amistre64):

10sin(c) - 5 sin(2c) = 20/pi 10sin(c) - 10 sin(c) cos(c) = 20/pi 10sin(c) (1- cos(c)) = 20/pi sin(c) - sin(c)cos(c) = 2/pi yeah, isnt really all that workable without some fancier stuff in my eyes

OpenStudy (rational):

then she wants to find the value of x at which f(x) achieves this average value

OpenStudy (amistre64):

can we do a newton rhapsody on f?

OpenStudy (dan815):

10*sin(x) -5*sin(2x) = 20/pi

OpenStudy (dan815):